# Distribution Prompting: Understanding the Expressivity of Language Models Through the Next-Token Distributions They Can Produce

### v1

@@ 5.3 Approximation of LM-generated Distributions
Building on the findings of Morris et al. (2024), who demonstrate that input text can be largely reconstructed from output distribution logits using a trained inversion model, we hypothesize that LM–generated distributions are intrinsically easier to approximate than distributions we experimented in § 5.1, irrespective of their entropy levels.
To evaluate this hypothesis, we conduct a comparative analysis between vanilla target distributions (Eq. 1) and target distributions derived from LM outputs. We conducted our experiments on Llama3.2-1B. We construct target distributions by randomly sampling a prompt—i.e., a sequence of random token indices—and using the model’s next-token output as the target distribution for evaluation. We compute the entropy of each LM-generated output distribution and use it as the target entropy in our training objective, as defined in Eq. 1. For each, we construct a corresponding vanilla next-token distribution with matched entropy. To mitigate initialization variance, we tune 25 prompt initializations per target and report the minimum loss. Hard prompt tuning is evaluated on 100 such distributions, each paired with a matched-entropy vanilla distribution.
We present the results in Fig. 8. As shown in Fig. 8(a), we fit a regression line over the minimum losses and their corresponding entropies across initializations. Hard prompt tuning on LM-generated distributions consistently yields lower loss than on vanilla ones, with a low R² value indicating weak correlation between entropy and loss. Fig. 8(b) reports the entropy of the next-token distributions produced by tuned hard prompts for both types of targets. The approximated entropy closely aligns with the target for LM-generated distributions, whereas vanilla targets show larger deviations.
@@ 5.4 Approximation of Variations of LM-Generated Distributions
To assess whether LM-generated distributions are not only easier to approximate, as demonstrated in § 5.3, but also robust under distributional variations, we extend our analysis to include shuffled LM-generated distributions and distributions derived from the outputs of other models.
@@
The analysis begins by evaluating the sensitivity of approximation performance to token assignments in the target distribution. We randomly shuffle the token indices of the LM-generated distributions from § 5.3, preserving their original probabilities. We refer to the resulting distributions as shuffled variants. Each original distribution is shuffled five times to reduce randomness, yielding five distinct variants for evaluation.
Following the previous setup, soft prompt tuning is performed on each distribution using five distinct initializations, and the minimum loss across runs is reported. Fig. 9 shows the minimum loss across shuffled variants, alongside those for the original LM-generated and matched vanilla distributions. Notably, even after shuffling, shuffle variants yield lower approximation loss than vanilla ones—especially in the moderate entropy regime where approximation is more difficult, as shown in Fig. 5.
@@
We further examine the transferability of LM-generated distributions by testing whether output probabilities can be effectively approximated across models. A source LM generates target distributions using the method in § 5.3, while a tuning LM performs prompt tuning on these distributions. The target distribution for the tuning LM is constructed by mapping shared tokens from the source LM and renormalizing their probabilities to ensure a valid distribution.
Soft prompt tuning is used to approximate the migrated distributions from the source models, with Llama3.2-1B as the tuning LM. Results in Fig. 10 show that, despite not being generated by the tuning model, the migrated distributions achieve lower approximation loss than matched-entropy vanilla distributions. This suggests that, despite differences in tokenization and architecture, LM outputs share underlying properties that support more effective approximation.
@@ 6 Conclusion and Discussion
In this work, we systematically study the expressiveness of LMs in terms of approximating target next-token probability distributions with the prompt tuning frameworks. We first analyze the expressiveness of different prompt tuning methods and fine that using continuous soft prompt tuning exhibits the best performance in distribution approximation. Through an information-theoretic lens, we design distributions of different entropy as tuning objectives and find that LMs had trouble prompting out distributions with moderate entropy values. Notably, we observe that having outlier tokens that are rare but receive distinctively higher probability values also reduces the difficulty for LMs to approximate. However, despite having moderate entropy value, distributions output by LMs themselves, are generally easier to approximate even under the less capable hard prompt tuning framework. Such distributions are still relatively easy to approximate if we add intervention, such as perturbation, to them. These general properties about the expressiveness of LMs can be a cornerstone for future research into more fine-grained studies.
It is natural to hypothesize that the difficulty for LMs to approximate moderate entropy distributions comes from the LM training objectives—LMs are trained to optimize the cross-entropy loss between a one-hot target distribution, which represents the single next token, and the predicted one, and such training objectives encourages the model to produce low-entropy distributions. However, our experiments have rejected this hypothesis—both randomly initialized and pretrained models exhibit the similar phenomenon. Additionally, if the above hypothesis is true, the approximation loss should be a monotonically increasing function of the target entropy, which is not the case according to our results. We hypothesize that such particular difficulty for approximating moderate-entropy distributions may come from either the Transformer (Vaswani et al., 2017) architecture or the softmax operator in the decoding stage. We leave the investigation of the root cause of this phenomenon for future work.
While language models are trained to predict the next token, the probabilistic nature of the output distribution makes it naturally a distribution proposer, from the perspectives of subjective randomness Bigelow et al. (2023) and computational creativity Peeperkorn et al. (2024), among others. Our work provides a first step to systematically analyze and understand the next-token distribution by LMs, and we hope our work can inspire more research in this direction.
@@
Our study could be extended in the following ways. First, we depend on the investigation of different target probability distributions entirely on the perspective of entropy, which is still a minor step compared to the whole picture of the understanding of the expressiveness of LMs. Second, we noticed the initialization of the prompt tuning actually matters in approximating certain distributions, which is also worthy of research. Finally, we still lack a clear understanding of the underlying cause of the phenomenon observed in this study, and further research is required to investigate its root cause. App. C presents a case study on instruction-tuned variants and POS-specific distributions. We hope this work motivates deeper exploration of LM distributional behavior and guides the design of more robust, interpretable models.
@@ Appendix A Entropy
We now establish the detailed calculation for value $m$ defined as the upper boundary for probability mass concentrate on outlier tokens in Eq. 3. Because of the linearity of entropy $H(p)$, we take the probability distribution of outlier tokens and the rest into separate consideration. For $k$ outlier tokens, the maximum entropy is reached when there is uniform distribution for them. As we assume that $m$ probability mass has fallen on these tokens, the maximum entropy for outlier tokens is $-m\log(\frac{m}{k})$. Similarly, for the rest tokens, the maximum entropy is $-(1-m)\log(\frac{1-m}{n-k})$. So in case we desire to find a probability distribution with the existence of $k$ outliers to meet a specific entropy $e$, we must satisfy the inequality:
$-m\log(\frac{m}{k})-(1-m)\log(\frac{1-m}{n-k})\geq e.$
@@ B.1 Prefix Initialization
We observed that the approximation loss varies across different initializations, as illustrated in Fig. 4. In the main body of this study, we report the minimum approximation loss obtained from five different initializations. While the absolute approximation values differ depending on the initialization, we find that the overall trends observed across the studied distributions remain consistent.
@@ B.2 Prefix Length
A longer prefix introduces more learnable parameters, which may potentially improve performance in approximating target distributions. To investigate this, we conducted an ablation study on prefix length using the vanilla distributions, as shown in § 5.1. We found that increasing the number of tokens in the prefix—thus making it more expressive—does not alter the general trend. The results of approximating distributions with varying entropy under different prefix length configurations are presented in Fig. 11.
@@ Appendix C Future Directions
We provide a case study investigating how well LMs approximate uniform distributions over tokens sharing the same part of speech. This illustrative example demonstrates the applicability of our proposed framework (Fig. 2) for analyzing the expressiveness of LMs. We envision that future work can further build on this framework to explore model behaviors from new interpretability perspectives.
We design a simple experiment to investigate the difficulty of approximating uniform distributions over tokens belonging to different parts of speech (POS). Specifically, we select three LMs from two distinct families and their instruction-tuned variants: Llama3.2-1B (Dubey et al., 2024), Qwen2.5-0.5B, and Qwen2.5-1.5B (Yang et al., 2024). To identify the POS tags of tokens from the model vocabularies, we utilize ScispaCy (Neumann et al., 2019) for tokenization and tagging. For simplicity, we restrict our analysis to tokens that are assigned a single, unambiguous POS tag. Fig. 12 shows the distribution of the eight most common POS tags across English tokens in the vocabularies of the two model families.
@@
We construct POS-wise uniform distributions by randomly sampling 100 tokens for nouns, proper nouns, verbs, and adjectives for 100 times, and create a uniform distribution each time. Following the same setup in App. B, we report the minimum approximation loss across 5 different prompt initializations.
@@
The experimental results are presented in Fig. 13. We observe that across all three comparisons, the instruction-tuned variants consistently exhibit higher approximation loss than their base counterparts. This suggests the presence of systematic distributional shifts introduced by instruction tuning. Additionally, we find that all evaluated LMs achieve lower approximation loss on adjectives compared to other POS.

### v2

@@ 5.3 Approximation of LM-generated Distributions
Building on the findings of Morris et al. (2024), who demonstrate that input text can be largely reconstructed from output distribution logits using a trained inversion model, we hypothesize that LM-generated distributions are intrinsically easier to approximate than distributions we experimented with in §˜5.1, irrespective of their entropy levels.
To evaluate this hypothesis, we conduct a comparative analysis between vanilla target distributions (Eq.˜1) and target distributions derived from LM outputs. We conducted our experiments on Llama3.2-1B. We construct target distributions by randomly sampling a prompt—i.e., a sequence of random token indices—and using the model’s next-token output as the target distribution for evaluation. We compute the entropy of each LM-generated output distribution and use it as the target entropy in our training objective, as defined in Eq.˜1. For each, we construct a corresponding vanilla next-token distribution with matched entropy. To mitigate initialization variance, we tune 25 prompt initializations per target and report the minimum loss. Hard prompt tuning is evaluated on 100 such distributions, each paired with a matched-entropy vanilla distribution.
We present the results in Fig.˜8. As shown in Fig.˜8(a), we fit a regression line over the minimum losses and their corresponding entropies across initializations. Hard prompt tuning on LM-generated distributions consistently yields lower loss than on vanilla ones, with a low R² value indicating weak correlation between entropy and loss. Fig.˜8(b) reports the entropy of the next-token distributions produced by tuned hard prompts for both types of targets. The approximated entropy closely aligns with the target for LM-generated distributions, whereas vanilla targets show larger deviations.
@@ 5.4 Approximation of Variations of LM-Generated Distributions
To assess whether LM-generated distributions are not only easier to approximate, as demonstrated in §˜5.3, but also robust under distributional variations, we extend our analysis to include shuffled LM-generated distributions and distributions derived from the outputs of other models.
@@
We begin by evaluating how sensitive approximation performance is to token assignments in the target distribution. To do so, we randomly shuffle token indices in LM-generated distributions from §˜5.3, preserving their original probabilities. These are referred to as shuffled variants, with each distribution shuffled five times to reduce randomness.
Following the prior setup, soft prompt tuning is applied using five random initializations, and the minimum loss is reported. Fig.˜9 shows the results for shuffled variants alongside original and matched vanilla distributions. Notably, shuffled variants still outperform vanilla ones—especially in the moderate entropy regime where approximation is harder (Fig.˜5).
@@
We further examine the transferability of LM-generated distributions by testing whether output probabilities can be effectively approximated across models. A source LM generates target distributions using the method in §˜5.3, while a tuning LM performs prompt tuning on these distributions. The target distribution for the tuning LM is constructed by mapping shared tokens from the source LM and renormalizing their probabilities to ensure a valid distribution.
Soft prompt tuning is used to approximate the migrated distributions from the source models, with Llama3.2-1B as the tuning LM. Results in Fig.˜10 show that, despite not being generated by the tuning model, the migrated distributions achieve lower approximation loss than matched-entropy vanilla distributions. This suggests that, despite differences in tokenization and architecture, LM outputs share underlying properties that support more effective approximation.
@@ 6 Conclusion and Discussion
In this work, we systematically study the expressiveness of LMs in approximating target next-token probability distributions under prompt tuning frameworks. We first analyze various prompt tuning methods and find that continuous soft prompts yield the best approximation performance. From an information-theoretic perspective, we design target distributions with varying entropy and observe that LMs struggle most with those of moderate entropy. Interestingly, the presence of outlier tokens—rare tokens with notably high probabilities—makes approximation easier. Moreover, distributions generated by LMs themselves are generally easier to approximate, even under hard prompt tuning, despite having moderate entropy. These distributions also remain relatively easy to approximate when perturbed. Overall, our findings reveal general properties of LM expressiveness that can inform more fine-grained future studies.
A natural hypothesis is that the difficulty LMs face in approximating moderate-entropy distributions stems from their training objective: minimizing cross-entropy with a one-hot target encourages low-entropy outputs. However, our experiments reject this explanation—both randomly initialized and pretrained models exhibit similar behavior. Moreover, if the hypothesis were correct, approximation loss would increase monotonically with target entropy, which our results do not support. We hypothesize that such particular difficulty for approximating moderate-entropy distributions may come from either the Transformer (Vaswani et al., 2017) architecture or the softmax operator in the decoding stage. We leave the investigation of the root cause of this phenomenon for future work.
While language models are trained to predict the next token, the probabilistic nature of the output distribution makes it naturally a distribution proposer, from the perspectives of subjective randomness Bigelow et al. (2023) and computational creativity Peeperkorn et al. (2024), among others. Our findings also have implications for interpretability and control: understanding which distributions are harder or easier to elicit can guide robust prompt design, distribution-aware decoding, and fine-grained behavioral analysis (e.g., POS-specific in App.˜C). Beyond prompting, such insights may support failure detection, model evaluation, and the development of decoding strategies that explicitly account for the reachable distribution space. Our work provides a first step to systematically analyze and understand the next-token distribution by LMs, and we hope this perspective spurs further exploration of LM expressivity and its applications.
@@
Our study could be extended in the following ways. First, we depend on the investigation of different target probability distributions entirely from the perspective of entropy, which is still a minor step compared to the whole picture of understanding the expressiveness of LMs. Second, we notice that the initialization of the prompt tuning actually matters in approximating certain distributions, which is also worthy of research. Ultimately, the root cause of the observed phenomenon remains unclear, warranting further investigation to better understand its underlying mechanisms. App.˜C presents a case study on instruction-tuned variants and POS-specific distributions. We hope this work motivates a deeper exploration of LM distributional behavior and guides the design of interpretable models.
Acknowledgements
We thank Ziqiao Ma and Jiayuan Mao for constructive discussions at the early stage of this work, as well as the anonymous reviewers and area chairs for the valuable feedback. This work is supported in part by NSERC RGPIN-2024-04395 and a Canada CIFAR AI Chair award to FS.
@@ Appendix A Entropy
We now establish the detailed calculation for value $m$ defined as the upper boundary for probability mass concentrate on outlier tokens in Eq.˜3. Because of the linearity of entropy $H(p)$, we take the probability distribution of outlier tokens and the rest into separate consideration. For $k$ outlier tokens, the maximum entropy is reached when there is uniform distribution for them. As we assume that $m$ probability mass has fallen on these tokens, the maximum entropy for outlier tokens is $m\log(\frac{m}{k})$. Similarly, for the rest tokens, the maximum entropy is $(1-m)\log(\frac{1-m}{n-k})$. So in case we desire to find a probability distribution with the existence of $k$ outliers to meet a specific entropy $e$, we must satisfy the inequality:
$m\log(\frac{m}{k})-(1-m)\log(\frac{1-m}{n-k})\geq e.$
@@ B.1 Prefix Initialization
We observed that the approximation loss varies across different initializations, as illustrated in Fig.˜4. In the main body of this study, we report the minimum approximation loss obtained from five different initializations. While the absolute approximation values differ depending on the initialization, we find that the overall trends observed across the studied distributions remain consistent.
@@ B.2 Prefix Length
A longer prefix introduces more learnable parameters, which may potentially improve performance in approximating target distributions. To investigate this, we conducted an ablation study on prefix length using the vanilla distributions, as shown in §˜5.1. We found that increasing the number of tokens in the prefix—thus making it more expressive—does not alter the general trend. The results of approximating distributions with varying entropy under different prefix length configurations are presented in Fig.˜11.
@@ Appendix C Future Directions
We provide a case study investigating how well LMs approximate uniform distributions over tokens sharing the same part of speech. This illustrative example demonstrates the applicability of our proposed framework (Fig.˜2) for analyzing the expressiveness of LMs. We envision that future work can further build on this framework to explore model behaviors from new interpretability perspectives.
We design a simple experiment to investigate the difficulty of approximating uniform distributions over tokens belonging to different parts of speech (POS). Specifically, we select three LMs from two distinct families and their instruction-tuned variants: Llama3.2-1B (Dubey et al., 2024), Qwen2.5-0.5B, and Qwen2.5-1.5B (Yang et al., 2024). To identify the POS tags of tokens from the model vocabularies, we utilize ScispaCy (Neumann et al., 2019) for tokenization and tagging. For simplicity, we restrict our analysis to tokens that are assigned a single, unambiguous POS tag. Fig.˜12 shows the distribution of the eight most common POS tags across English tokens in the vocabularies of the two model families.
@@
We construct POS-wise uniform distributions by randomly sampling 100 tokens for nouns, proper nouns, verbs, and adjectives for 100 times, and create a uniform distribution each time. Following the same setup in App.˜B, we report the minimum approximation loss across 5 different prompt initializations.
@@
The experimental results are presented in Fig.˜13. We observe that across all three comparisons, the instruction-tuned variants consistently exhibit higher approximation loss than their base counterparts. This suggests the presence of systematic distributional shifts introduced by instruction tuning. Additionally, we find that all evaluated LMs achieve lower approximation loss on adjectives compared to other POS.